8. Inputs and outputs Clancys Performance Pizza is a small restaurant in Houston that sells gluten-free pizzas. Clancys verFill in the blanks to complete the Marginal Product of Labor column for each worker. Labor (Number of workers) Output (PizzasProduction Function QUANTITY OF OUTPUT (Pizzas) LABOR (Number of workers) Suppose that labor is Clancys only variable cost aUse the orange points (square symbol) to plot Clancys total cost curve on the following graph using the quantities from theTrue or False: The shape of the production function reflects the law of increasing marginal returns. O True O False
Answer #1

Answer:

In the short run, these workers are variable inputs, and the ovens are fixed inputs.

Labor

Output

Marginal Product of Labor

0

0

-

1

60

60-0=60

2

100

100-60=40

3

130

130-100=30

4

150

150-130=20

5

160

160-150=10

Production Function QUANTITY OF OUTPUT (Pizzas) LABOR (Number of workers)

Number of workers

output

Fixed cost ($)

Variable cost ($)

Total cost ($) = FC+VC

0

0

20

0

20

1

60

20

30

50

2

100

20

60

80

3

130

20

90

110

4

150

20

120

140

5

160

20

150

170

Total Cost TOTAL COST (Dollars) 0 20 180 200 40 60 80 100 120 140 160 QUANTITY OF OUTPUT (Pizzas)

False

Explanation:

The production function reflects the law of decreasing return, because MPL is decreasing.